OpenStudy (anonymous):

Is the derivative of y=2sec(3x) = 6sec^2(3x)*3tan(3x)

OpenStudy (anonymous):

I'll double-check, but it looks like there is too much going on there.

jimthompson5910 (jim_thompson5910):

The derivative of sec(x) is sec(x)*tan(x)

jimthompson5910 (jim_thompson5910):

so I'm not sure how you got the sec^2

OpenStudy (anonymous):

\(\large 2sec(3x)=\frac{2}{cos(3x)}\) Can use quotient rule (and a little chain rule), to get \(\large \frac{(-2)(3)(-sin(3x))}{cos^2(3x)}\)

OpenStudy (anonymous):

I combined the two sec(3x)?

jimthompson5910 (jim_thompson5910):

oh, when you derive 2*sec(x), you pull out the 2 and ignore it when you derive

OpenStudy (anonymous):

y=2sec(3x) (2)*(3sec(3x))+(0)*(sec(3x)) 2*3sec(3x)*3tan(3x)+0*sec(3x) 6sec(3x)*3tan(3x)+sec(3x) 6sec^2(3x)*3tan(3x) this is what i did.

jimthompson5910 (jim_thompson5910):

here's what you should do y = 2*sec(3x) y' = d/dx[2*sec(3x)] y' = 2*d/dx[sec(3x)]*d/dx[3x] y' = 2*sec(3x)*tan(3x)*3 y' = 6*sec(3x)*tan(3x)

OpenStudy (anonymous):

so i dont use the product rule?

jimthompson5910 (jim_thompson5910):

no

jimthompson5910 (jim_thompson5910):

2 is a constant which you can pull out

jimthompson5910 (jim_thompson5910):

leaving you with you to derive sec(3x) to get sec(3x)*tan(3x) then you multiply that with the derivative of 3x (the inner function), which is 3

OpenStudy (anonymous):

do i always do that when its multiplied by a constant?

jimthompson5910 (jim_thompson5910):

yes

jimthompson5910 (jim_thompson5910):

In general, If y = k*f(x) then y' = k*f ' (x) where k is a constant and f(x) is some function

OpenStudy (anonymous):

then why did you multiply by 3x at the end

OpenStudy (anonymous):

derivative of 3x*

jimthompson5910 (jim_thompson5910):

that's from the chain rule

OpenStudy (tkhunny):

(d/dx)sec(x) = sec(x) * tan(x) (d/dx)sec(3x) = 3 * sec(3x) * tan(3x) (d/dx)2sec(3x) = 6 * sec(3x) * tan(3x) Notice how adding the constant 2 doesn't effect anything. It just comes along for the ride at the end.

OpenStudy (anonymous):

ohhh okayy.. i got it!

jimthompson5910 (jim_thompson5910):

that's great

OpenStudy (anonymous):

thanks you guys!

jimthompson5910 (jim_thompson5910):

yw